OpenStudy (anonymous):

The dimensions shown are an approximation. Using these dimensions, what is the approximate area of triangle ABD? How many sq miles?

OpenStudy (anonymous):



OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Area of a triangle=1/2 x base x height Area=1/2 x 4.5 x 4 Area=2 x 4.5 Area=9

OpenStudy (bibby):

the base of ABD is not 4.5

TheSmartOne (thesmartone):

oh..

OpenStudy (anonymous):

how would that be split

OpenStudy (bibby):

what you might be able to do is \(\huge \frac{1}{2}*b*h=\frac{1}{2}*\frac{4.5}{2}*4=\frac{4.5*\cancel{4}}{\cancel{4}}\)

OpenStudy (anonymous):

so that leaves me with 4.5

OpenStudy (bibby):

yeah, probably, maybe